OpenStudy (anonymous):

if you were to have -3x + y = 1 , how would you go about doing that ?

OpenStudy (mertsj):

Depends on what I wanted to do with it. What do you want to do with it?

OpenStudy (anonymous):

Well , on the problem it shows a graph, but theres no way i can show it to you. but , here's the problem Isaac collected data for an experiment. The water level of his experiment increased at a constant rate. Isaac could model his data with the equation -3x + y = 1. Which is the graph of Isaac's data?

OpenStudy (mertsj):

Put it in slope-intercept form. That is put it in the form y = mx +b

OpenStudy (mertsj):

What do you get?

OpenStudy (mertsj):

rhonda?

OpenStudy (mertsj):

you there?

OpenStudy (anonymous):

Just a second .

OpenStudy (anonymous):

So , would you do down 3 and over 3 ?

OpenStudy (mertsj):

Did you write it in the y = mx+b form?

OpenStudy (anonymous):

Yes.

OpenStudy (mertsj):

And what is it?

OpenStudy (anonymous):

well , it's kind of hard for me to explain , but I got the answer wrong , and I understand now . thanks.

OpenStudy (mertsj):

It's y = 3x +1. That is a line whose y intercept is 1 and whose slope is 3/1. So choose the graph that crosses the y axis at 1 and from that point goes up 3 and 1 to the right.